OpenStudy (anonymous):

y=-x+6 7y=5x-42 solve for x and y

OpenStudy (lgbasallote):

see this tutorial on systems of linear equations http://openstudy.com/updates/4fa7aa68e4b029e9dc388281 tell me if you dont get something and ill explain it to you

OpenStudy (anonymous):

Notice how the second equation includes y, so replace y with the first equation: \[7(-x+6) = 5x-42\] Expanding: \[-7x + 42 = 5x - 42\] Then group terms: \[84 = 12x\] Divide to solve for x: \[x=7\] Now substitute x in any of the equations; the first is simpler, so let's start there. \[y=-(7)+6\] Therefore y=-1 We can check this by plugging in x and y in both equations and hope these values make each equation true. \[-1 = -7 + 6\] \[7(-1)=5(7)-42\] Both left and right sides are equal; therefore the solutions are x=7, y=-1.